Question 4: A gold mine extract gold from rock using toxic substances. For each additional tonne of rock the marginal benefit of extraction are MB =12-q where q is the quantity of rock. The extraction process causes marginal environmental damage equal to MD = 2 q. a) Without regulation, how much rock would the gold mine extract? b) With strict liability laws, how much gold would the mine extract?

Question 5: Consider the same scenario as in question 4. A) How high is the Pigovian tax? B) How much rock would the mine extract under that tax?

Question 6: What is the efficiency gain from implementing a Pigovian tax?
Answer #1

Answer 4: Without regulation MD = 0

Equilibrium is where MB = MD

12-q = 0

q = 12 units.

b) With strict liability laws,

MD = 2q will be borne by the producer.

hence equilibrium will be where

12-q = 2q

3q = 12

q = 4 units.

Answer 5 Pigouvian tax is $ 2 per unit.

After tax marginal benefit will be 12-q-2q = 12-3q

SO mine will extract rock equal to a point where

12-3q = 0

q = 4 units